# It Cannot Be Right If It Was Written by AI: On Lawyers’ Preferences of Documents Perceived as Authored by an LLM vs a Human

### v1

@@ 5 Discussion
Before we approach individual issues, we must address the overarching trend, which, while not the subject of our study, is noteworthy. The documents were prepared to be of comparable quality, terminology, and structure, established via cross-validation within the authors’ team and by approaching an experienced attorney outside the authors’ team. However, the Verbose document, on average, outperformed the Brief document in the evaluation, both in terms of correctness and language quality. Detailed investigation of the trend reveals though that, e.g., a Brief document labelled as Human-crafted outperforms Verbose documents labelled as AI-generated. However, the difference is not significant enough to overcome the general preference of the Verbose document by both lawyers and law students. Such a preference may be rooted in the tendency of legal professionals to associate length with precision.
@@
Additionally, it reflects the tacit nature of legal knowledge. Civil law, of which the acknowledgement of debt we used in the study is the prime example, uses dispositive norms. Unlike cogent norms, dispositive norms allow contracting parties to reach an agreement which is, in detail, different from the law. It might be that lawyers preferred the Verbose document in their expectation that it captures the will of the parties expressly and fully and does not leave any part of the contract tacitly bound to the Civil Code. While we noted the general preference for the Verbose document over the Brief one, the AI-generated Verbose document was rated worse than the Brief document labelled Human-crafted. Lawyers expect longer Human-crafted documents to be a sign of eloquence, while longer AI-generated documents indicate uncertainty and an inability to operate in the legal domain. The preference for Verbose documents should be investigated further in terms of underlying motivations.
@@
In terms of individual trends within the scope of this study, participants were more critical when scoring and commenting on language quality over correctness. Negative comments about language were significantly more prevalent for both Brief and Verbose documents when presented as AI-generated. The risks of LLMs, associated mainly with the prevalent hallucinations, are well-known among both lawyers and law students. As such, participants may have entered the survey expecting to see these issues manifest when dealing with AI-generated documents. Since both documents were prepared to meet legal requirements, and the acknowledgement of debt is straightforward, participants may have focused on language-related issues instead of correctness. Specialized legal jargon is not known for its brevity and conciseness. As a result, most participants can point out weaknesses based on individual stylistic preferences. Castelo et al [15] identified the aversion to algorithms as higher in tasks perceived as subjective.
@@
The correctness is objective, while language issues are affected by individual preferences, such as using specific phrases. Such an aversion to distinguishing between specific and objective tasks may be another explanation for negative comments directed at the documents’ language quality when presented as AI-generated. Our interpretation is further supported by the heightened occurrence of negative comments tied to more subjective aspects of correctness. For example, the ’formal correctness’ of the documents presented as AI-generated drew a lot of negative comments, as it is again a subjective category where individual preferences matter.
@@
In support of this interpretation, we noted that none of the documents, either Brief or Verbose, whether labelled as AI-generated or Human-crafted, received a negative comment regarding legal requirements and expected legal consequences. While individual preferences may influence other properties, these are the most objective, as they are directly based on the law. Lawyers may object to subjective properties, but it is difficult to dispute the ability of documents to give rise to specific legal consequences. The interpretation may serve as a stepping stone for other studies with a more nuanced approach towards evaluating documents’ properties. We focused on correctness and language quality, but both broadly defined categories contain subjective sub-categories. Mapping the impact of these variables would have a profound impact on the further use of LLMs within the legal domain.
@@
Documents labelled as Human-crafted were consistently rated better than AI-generated documents in terms of correctness. However, there is nuance to these results, which warrants further attention. Thematic analysis has shown lawyers criticizing the AI-generated Brief document in terms of its coherence and logical structure, and formal correctness. The prevalence of negative comments in these areas was lower in the AI-generated Verbose document. Inherent negative expectations regarding the language capabilities of LLMs may cause this. The legal jargon often contains complicated sentences with conditions and caveats. The lawyers seem to expect AI to be unable to use the language proficiently enough to structure a legal text and perceive the AI-generated Brief document as supportive evidence for their stance.
@@
On the other hand, and quite surprisingly, the AI-generated Brief document attracted more positive comments on legal requirements and expected legal consequences than its Human-crafted Brief counterpart. Well aware of risks related to hallucinations of factual information, lawyers were probably surprised by the factuality of the Brief Document presented as AI-generated. Lawyers are prepared to encounter imperfect or faulty AI-generated documents. As a result, they felt the urge to comment positively on what they otherwise perceived as ’normal’ or ’expected’ within human-crafted documents. The aspect of expectation was presented even in the Human-crafted Verbose document, receiving more negative comments in terms of coherence and logical structure than the AI-generated Verbose document. Lawyers seem to have low expectations of AI-generated documents. On the other hand, their opinion of Human-crafted documents may be affected by the projection of how the document would have looked should they have been the ones drafting it. Such a projection can carry with it a negative sentiment.
@@
Additionally, documents labelled as Human-crafted received better overall scores and more positive and less negative comments. These findings are in line with the existing literature on algorithmic aversion. While the previous research reported lower evaluation of AI-generated content in case of Airbnb profiles [40], emails [58], artworks [78], music [86], translations [2], news articles [93] or health prevention messages [57], we identified the lower evaluation of AI-generated legal content. Automated legal content generation may also lack the required level of personalized care about one’s legal issues. Such an outcome aligns with concerns about insufficient personalization of cases observed by Longoni et al [60] within the healthcare sector.
@@
Although documents labelled as AI-generated were consistently rated worse, and the thematic analysis yielded more negative comments on both their linguistic and legal quality, the results of the open-ended question about possible future automation reveal that practitioners certainly envisage full automation of (at least) simple legal documents such as the acknowledgement of debt in this experiment. While there is a noticeable scepticism towards AI-generated legal documents, lawyers consider full automation in the future almost certain. The interpretation holds regardless of the lower rating of documents labelled as AI-generated in both correctness and language quality.
@@ 6 Implications for Legal Practice
The disruptive potential of LLMs is immense. The most often appearing opportunities are decreasing costs associated with legal services and increasing access to justice. ChatGPT’s appeal partly lies in the democratization of legal services. Lawyers’ services are often perceived as prohibitively costly. Laymen already use LLMs to address their legal needs [32], and the trend is likely to grow.
@@
These two are communicating vessels. Inaccuracies, either real or perceived, can hamper the use of LLMs in the legal domain. As our study shows, the perception of AI-generated documents must be studied separately from their objective accuracy. The performance of LLMs in individual tasks, either low-level or complex, is just one part of the story. Even objectively accurate documents can be perceived negatively when their recipients believe the documents were AI-generated.
@@
However, even if appropriately used, the perception of lower-quality of AI-generated content harms the chances as well. Additionally, people from lower-income groups may not have to approach government agencies or courts with AI-generated content to be negatively affected. The mere perception that a person is using AI-generated content can lower their chances of addressing their legal needs. In a broader context, unequal viewing of similar legal documents with different origins (AI, human) may in turn contribute to increased social inequalities.
Our results have shown that the label of AI-generated content carries with it negative associations, such as broad indication of the lower quality or general inadequacy of a legal document.
As a result, mandatory disclosure of using LLMs may manifest as a conflict between fairness and transparency. Algorithmic aversion may become another bias, which will require practical measures to address its detrimental impact. The high-level requirements for fair and just trial may not be enough to address implicit biases, whether racial, gender, or AI-related. Instead of better access to justice aided by ChatGPT, LLMs, and AI in general, we may be left with constraints placed upon those needing legal assistance the most.
@@ 7 Limitations
Our study’s limited number of participants may have introduced some bias to the results. As a result, some of the reported results, particularly those related to ’correctness’ and ’language quality’, may not be statistically significant. This limited evaluation does not provide a sufficiently detailed look at the various properties of the documents. A more comprehensive thematic analysis is a necessity to address this limitation.
Our study evaluated the acknowledgement of debt documents. While the document is commonly used and is a relevant case study, our findings may not extend to other legal documents. Additionally, acknowledgement of debt is simple. Different document types have more complex legal requirements or linguistic and structural properties. Exploring a more comprehensive range of legal documents would lead to more robust data.
@@
Addressing these limitations in future research will be crucial to achieving more generalizable and potentially practical results.
@@ 8 Conclusions and Future Work
Our study shed light on how lawyers perceive AI-generated legal documents compared to Human-crafted documents. Participants rated the AI-generated documents worse than their Human-crafted counterparts.
Participants were more critical of language quality than correctness, with AI-generated documents receiving significantly more negative comments. The aversion to AI-generated legal content aligns with existing research on algorithmic aversion in other domains. Documents labelled as human-crafted received higher overall scores and more positive comments, reflecting a possible general preference for the involvement of human experts in producing legal content.
Despite the prevalent negative perception, our study reveals significant optimism among the participants about the future of automating legal content production. They highlighted factors such as the indistinguishability between Human-crafted and AI-generated content and the ability of AI-generated content to produce legal consequences, indicating a potential shift in perception in the future.
@@
While our findings are significant, they should be a starting point for further research. The potential repercussions for providing legal services, particularly for lower-income groups and the mandatory designation of AI-generated documents, require careful evaluation. Further research must focus on other populations, larger samples, and more complex documents, as well as a nuanced understanding of documents’ objective and subjective properties and the prior knowledge of intended recipients.
The perception of AI-generated documents still needs to be studied, both regarding the variables and their subsequent impact on legal practice.

### v2

@@ 5 Discussion
Before we approach individual issues, we must address the overarching trend, which, while not the subject of our study, is noteworthy. The documents were prepared to be of comparable quality, terminology, and structure, established via cross-validation within the authors’ team and by approaching an experienced attorney outside the authors’ team. However, the Verbose document, on average, outperformed the Brief document in the evaluation, both in terms of correctness and language quality. A detailed investigation of the trend reveals that the Brief document labeled as human-crafted outperforms the Verbose document labeled as AI-generated. However, the difference is not significant enough to overcome lawyers’ and law students’ general preference for the Verbose document. Such a preference may be rooted in the tendency of legal professionals to associate length with precision.
@@
Additionally, it reflects the implicit nature of legal knowledge. Civil law, which the acknowledgement of debt we used in the study is the prime example, uses dispositive norms. Unlike cogent norms, dispositive norms allow contracting parties to reach an agreement which is, in detail, different from what is stipulated in the respective Act. It might be that lawyers preferred the Verbose document in their expectation that it captures the will of the parties expressly and fully and does not leave any part of the contract tacitly bound to the Civil Code. While we noted the general preference for the Verbose document over the Brief one, the AI-generated Verbose document was rated worse than the Brief document labeled human-crafted. Lawyers expect longer human-crafted documents to be a sign of eloquence, while longer AI-generated documents indicate uncertainty and an inability to operate in the legal domain. The preference for Verbose documents should be investigated further regarding underlying motivations.
@@
In terms of individual trends within the scope of this study, participants were more critical when scoring and commenting on language quality over correctness. Negative comments about language were more prevalent for both Brief and Verbose documents when presented as AI-generated. The risks of LLMs, associated mainly with the prevalent hallucinations, are well-known among both lawyers and law students. As such, participants may have entered the survey expecting to see these issues manifest when dealing with AI-generated documents. Since both documents were prepared to meet legal requirements, and the acknowledgement of debt is straightforward, participants may have focused on language-related issues instead of correctness. Specialized legal jargon is not known for its brevity and conciseness. As a result, most participants can point out weaknesses based on individual stylistic preferences. Castelo et al [15] identified the aversion to algorithms as higher in tasks perceived as subjective.
@@
The correctness is objective, while language issues are affected by individual preferences, such as using specific expressions. Such an aversion to distinguishing between specific and objective tasks may be another explanation for negative comments directed at the documents’ language quality when presented as AI-generated. Our interpretation is further supported by the heightened occurrence of negative comments tied to more subjective aspects of correctness. For example, the ’formal correctness’ of the documents presented as AI-generated drew a lot of negative comments, as it is again a subjective category where individual preferences matter.
@@
In support of this interpretation, we noted that none of the documents, either Brief or Verbose, whether labeled as AI-generated or human-crafted, received a negative comment regarding legal requirements and expected legal consequences. While individual preferences may influence other properties, these are the most objective, as they are directly based on the law. Lawyers may object to subjective properties, but it is difficult to dispute the ability of documents to give rise to specific legal consequences. The interpretation may serve as a stepping stone for other studies with a more nuanced approach towards evaluating documents’ properties. We focused on correctness and language quality, but both broadly defined categories contain subjective sub-categories. Mapping the impact of these variables would have a profound impact on the further use of LLMs within the legal domain.
@@
Documents labeled as human-crafted were consistently rated better than AI-generated documents in terms of correctness. However, there is nuance to these results, which warrants further attention. Thematic analysis has shown lawyers criticizing the AI-generated Brief document in terms of its coherence, logical structure, and formal correctness. The prevalence of negative comments in these areas was lower in the AI-generated Verbose document. Inherent negative expectations regarding the language capabilities of LLMs may cause this. The legal jargon often contains complicated sentences with conditions and caveats. The lawyers expect AI to be unable to use the language proficiently enough to structure a legal text and perceive the AI-generated Brief document as supportive evidence for their stance.
@@
On the other hand, quite surprisingly, the Brief document labeled as AI-generated attracted more positive comments on legal requirements and expected legal consequences than its Brief counterpart labeled human-crafted. Well aware of risks related to hallucinations of factual information, lawyers were probably surprised by the factuality of the Brief Document presented as AI-generated. Lawyers are prepared to encounter imperfect or faulty AI-generated documents. As a result, they felt the urge to comment positively on what they otherwise perceived as ’normal’ or ’expected’ within human-crafted documents. The aspect of expectation was presented even in the human-crafted Verbose document, receiving more negative comments regarding coherence and logical structure than the AI-generated Verbose document. Lawyers seem to have low expectations of AI-generated documents. On the other hand, their opinion of human-crafted documents may be affected by the projection of how the document would have looked should they have drafted it. Such a projection can carry with it a negative sentiment.
@@
Additionally, documents labeled as human-crafted received better overall scores and more positive and less negative comments. These findings are in line with the existing literature on algorithmic aversion. While the previous research reported lower evaluation of AI-generated content in case of Airbnb profiles [42], emails [60], artworks [81], music [89], translations [2], news articles [97] or health prevention messages [59], we identified the lower evaluation of AI-generated legal content. Automated legal content generation may also lack the required level of personalized care about one’s legal issues. The outcome aligns with concerns in the healthcare sector as Longoni et al [62] mention lack of perceived personalization as one source of resistance to the deployment of artificial intelligence.
@@
Although documents labeled as AI-generated were consistently rated worse, and the thematic analysis yielded more negative comments on both their linguistic and legal quality, the results of the open-ended question about possible future automation reveal that practitioners certainly envisage full automation of (at least) simple legal documents such as the acknowledgement of debt. While there is a noticeable scepticism towards AI-generated legal documents, lawyers consider full automation in the future almost certain. The interpretation holds regardless of the lower rating of documents labeled as AI-generated in both correctness and language quality.
@@ 6 Implications for Legal Practice
The disruptive potential of LLMs is immense. The most often appearing opportunities are decreasing costs associated with legal services and increasing access to justice. ChatGPT’s appeal partly lies in the democratization of legal services. Lawyers’ services are often perceived as prohibitively costly. Laypeople already use LLMs to address their legal needs [34], and the trend is likely to grow.
@@
These two are communicating vessels. Inaccuracies, either real or perceived, can hamper the use of LLMs in the legal domain. Our study shows that the perception of AI-generated documents must be studied separately from their objective performance. The performance of LLMs in individual tasks, either low-level or complex, is just one part of the story. Even objectively accurate documents can be perceived negatively when their recipients believe they are AI-generated.
@@
However, even if appropriately used, the perception of lower-quality of AI-generated content also harms their chances. Additionally, people from lower-income groups may not have to approach government agencies or courts with AI-generated content to be negatively affected. The mere perception that a person is using AI-generated content can lower their chances of addressing their legal needs. In a broader context, unequal viewing of similar legal documents with different origins (AI, human) may, in turn, contribute to increased social inequalities.
Our results have shown that the label of AI-generated content carries negative associations, such as broad indication of the document’s lower quality or general inadequacy to cause legal consequences.
As a result, mandatory disclosure of using LLMs may manifest as a conflict between fairness and transparency. Algorithmic aversion may become another bias, which will require practical measures to address its detrimental impact. The high-level requirements for fair and just trial may not be enough to address implicit biases, whether racial, gender, or AI-related. Instead of better access to justice aided by ChatGPT, LLMs, and AI in general, we may be left with implicit constraints placed upon those needing legal assistance the most.
@@ 7 Limitations
Our study’s limited number of participants (n=75) may have introduced bias to the results. As a result, some of the reported results are not statistically significant. Furthermore, the evaluation of the correctness and language quality does not sufficiently capture the various properties of the documents. A more comprehensive analysis is necessary to address this limitation.
Our study evaluated the acknowledgement of debt. While the document is commonly used and is a relevant case study, our findings may not extend to other legal documents. Additionally, acknowledgement of debt is simple. Different document types have more complex legal requirements or linguistic and structural properties. Exploring a more comprehensive range of legal documents would lead to more robust data.
@@
Addressing these limitations in future research will be crucial in achieving more generalizable and potentially practical results.
@@ 8 Conclusions and Future Work
Our study shed light on how lawyers perceive legal documents labeled as AI-generated compared to documents labeled as human-crafted. Participants rated documents labeled as AI-generated documents worse than their counterparts presented as human-crafted.
Participants were more critical of language quality than correctness, with AI-generated documents receiving more negative comments. The aversion to AI-generated legal content aligns with existing research on algorithmic aversion in other domains. Documents labeled as human-crafted received higher overall scores and more positive comments, reflecting a possible general preference for the involvement of human experts in producing legal content.
Despite the prevalent negative perception, our study reveals significant optimism among the participants about the future of automated legal content production. They highlighted factors such as the indistinguishability between human-crafted and AI-generated content and the ability of AI-generated content to produce legal consequences, indicating a potential shift in perception in the future.
@@
While our findings are significant, they should be a starting point for further research. The potential repercussions of providing legal services—particularly for lower-income groups—and the mandatory labeling of AI-generated documents require careful evaluation. Further research must focus on other populations, larger samples, and more complex documents, as well as a nuanced understanding of documents’ objective and subjective properties and the prior knowledge of intended recipients.
The perception of AI-generated documents still needs to be studied regarding the variables and their subsequent impact on legal practice.
[…]
Acknowledgements We thank Bettina Bacher, and the two anonymous reviewers, for their helpful comments on previous versions of this paper. Work of Jakub Harasta and Tereza Novotná was supported by the Grant Agency of Masaryk University (GAMU) project “Forensic Support for Building Trust in Smart Software Ecosystems” (no. MUNI/G/1142/2022).